# Debiasing Watermarks for Large Language Models via Maximal Coupling

### v1

@@ A.5.1 Additional Simulation Studies
A.5.2 Further Implementation Details on Language Model Experiments
@@
A.5.3 Ablation Study on the Parameters
@@
Table 5: Comparison of the S-BERT similarity scores and the true positive rates (TPR) among different watermarking schemes. The significance level is set to $\alpha=0.01$. The higher the S-BERT similarity score, the less the distortion. The higher the TPR, the better the detection power. The higher the TPR aug., the more robust the watermarking scheme is against the substitution attack.
Table 5 present the analogous results to Table 1 for the Llama-3.2-1B-Instruct model. The results are consistent with the main results, showing that our proposed watermarking scheme outperforms the Gumbel-max method in terms of the text distortion, and more competitive TPR than the green/red list and DiPMark methods.
@@
Table 6 and Table 7 present the analogous results with a wider range of parameters. Table 7 contains the results for the Llama model, while Table 6 contains the results for the Phi model. From both tables, we observe a positive correlation between the distortion of the generation (lower S-BERT scores) and the detection power (higher TPR) for the green/red list method, echoing existing theory and empirical findings (Kirchenbauer et al., 2023a; Fernandez et al., 2023; Piet et al., 2023).
@@
Table 8 presents the analogous results to Table 2 for the Llama model. Here, we observe that for the Llama model, the Gumbel-max method has a higher rate of repeated tokens compared to all other methods.
@@
Table 9 presents the analogous results to Table 3 for the Llama model. The results show that our proposed watermarking scheme has a lower average rejection rate compared to the Gumbel-max method, and this is consistent with the main results.
@@
Table 10: Same as Table 4, but for the Llama model.
Table 10 presents the analogous results to Table 4 for the Llama model. The results show higher true positive rate for the sum test compared to the higher criticism when the number of tokens used to detect the watermark is relatively small, and this consistent with the main results.
@@
A.5.4 A Single Green List
Notice that our theory requires that the green lists ${\mathcal{G}}$ and the uniform random variables $\zeta$ be independent, and our scheme does not necessitate creating a new green list before generating each token. Instead, as long as the user does not have access to the watermark key, a priori, we can simply generate a single green list at the beginning and use it for generating all the tokens. This way, we only need to share a single pseudorandom function to generate random variable $\zeta$’s. We performed an ablation study on this in Table 11. In particular, we compare the results of our watermarking scheme when we use a single green list and when we use disparate green lists. We report the BERT similarity scores and the TPR of the test statistic based on the sum of the $\zeta$’s and the higher criticism statistic (in the form TPR for sum/TPR for higher criticism). We conclude that having a single green list does not significantly impact our watermarking scheme when a larger context window is used to generate $\zeta$. An analogous observation was also made in Zhao et al. (2023) for the green/red list soft watermarking scheme. However, it is unclear whether having the same green list for all the tokens will allow an adversary to have more power to learn the exact green list generator and produce attacks that can fool the detector, which is an interesting direction for future work.

### v2

@@ A.6.1 Additional Simulation Studies
A.6.2 Further Implementation Details on Language Model Experiments
@@
A.6.3 Ablation Study on the Parameters
Table 5: Comparison of the S-BERT similarity scores and the true positive rates (TPR) among four watermarking schemes: (1) the Gumbel-max watermark (Aaronson and Kirchner, 2023), (2) the green/red list watermark (Kirchenbauer et al., 2023a), (3) the DiPmark method Wu et al. (2023), and (4) our proposed one.
@@ A.6.3 Ablation Study on the Parameters
TPR para. 1.0000 0.4000 0.7800 0.9200
Supplementary Table 5 present the analogous results to Table 1 for the Llama-3.2-1B-Instruct model. The results are consistent with the main results, showing that our proposed watermarking scheme outperforms the Gumbel-max method in terms of the text distortion, and more competitive TPR than the green/red list and DiPMark methods.
@@
Table 7: Same as Supplementary Table 6, but for the Llama model.
@@
Supplementary Table 6 and Supplementary Table 7 present the analogous results with a wider range of parameters. Supplementary Table 7 contains the results for the Llama model, while Supplementary Table 6 contains the results for the Phi model. From both tables, we observe a positive correlation between the distortion of the generation (lower S-BERT scores) and the detection power (higher TPR) for the green/red list method, echoing existing theory and empirical findings (Kirchenbauer et al., 2023a; Fernandez et al., 2023; Piet et al., 2023).
@@
Supplementary Table 8 presents the analogous results to Table 2 for the Llama model. Here, we observe that for the Llama model, the Gumbel-max method has a higher rate of repeated tokens compared to all other methods.
@@
Supplementary Table 9 presents the analogous results to Table 3 for the Llama model. The results show that our proposed watermarking scheme has a lower average rejection rate compared to the Gumbel-max method, and this is consistent with the main results.
@@
Supplementary Table 10 presents the analogous results to Table 4 for the Llama model. The results show higher true positive rate for the sum test compared to the higher criticism when the number of tokens used to detect the watermark is relatively small, and this consistent with the main results.
@@
A.6.4 A Single Green List
Notice that our theory requires that the green lists ${\mathcal{G}}$ and the uniform random variables $\zeta$ be independent, and our scheme does not necessitate creating a new green list before generating each token. Instead, as long as the user does not have access to the watermark key, a priori, we can simply generate a single green list at the beginning and use it for generating all the tokens. This way, we only need to share a single pseudorandom function to generate random variable $\zeta$’s. We performed an ablation study on this in Supplementary Table 11. In particular, we compare the results of our watermarking scheme when we use a single green list and when we use disparate green lists. We report the BERT similarity scores and the TPR of the test statistic based on the sum of the $\zeta$’s and the higher criticism statistic (in the form TPR for sum/TPR for higher criticism). We conclude that having a single green list does not significantly impact our watermarking scheme when a larger context window is used to generate $\zeta$. An analogous observation was also made in Zhao et al. (2023) for the green/red list soft watermarking scheme. However, it is unclear whether having the same green list for all the tokens will allow an adversary to have more power to learn the exact green list generator and produce attacks that can fool the detector, which is an interesting direction for future work.